Mr. McTier invests $600 at 9% per year, and $800 at 12% per year. During what time period will the $1,400 earn between $525 and $975, inclusive, in simple interest?
Answer by mananth(16949)

$600 at 9% per year,
Interest for 1 year = 0.09*600= $54
and $800 at 12% per year.
Interest for 1 year = 0.12*800= $ 96
Total for 1 year = $150
..
$150 1 year
For $525 --- 525/150 = 3.5 years
..
975 /150 = 6.5 years
..
Time frame is 3.5 years to 6.5 years